10 Using the bond enthalpies provided, calculate the heat of reaction, ΔH, for:

C2H4(g) + H2(g) → C2H6(g) 

Given the following table of bond enthalpies:

C — H

413 kJ


C = C

614 kJ


C —C

348 kJ


H — H

436 kJ



1
Expert's answer
2021-02-04T07:06:45-0500


C2H4(g) + H2(g) → C2H6(g) 


C2H4 molecule: 1 C=C (624 kJ) + 4 C-H (4 x 413 = 1652 kJ)

H2 molecule: 1 H-H (436 kJ)

Totally on the right side of the equation:

624 + 1652 + 436 = 2712 (kJ)

C2H6 molecule: 1 C-C (348 kJ) + 6 C-H (6 x 413 = 2478 kJ)

Totally on the left side of the equation:

348 + 2478 = 2826 (kJ)

Bonds broken – bonds formed: 2712 – 2826 = -114 (kJ)